Using CEA2, calculate the equilibrium composition of a mixture of 1 mol of CO2 , 2 mol of H2O, and 7.52 mol of N2 at 1 bar and temperatures ranging from 500 to 3000 K in steps of 100 K. Then plot the mole fraction of NO as a function of temperature.


1
Expert's answer
2021-08-06T04:30:45-0400

CO2(g)+H2(g)⇔CO(g)+H2(g)1100moles at t=0(1-x)(1-x)xxmoles at equilibrium

Given that volume %

% of CO=0.16

The moles at equilibrium "=1\u2212x+1\u2212x+x+x=2"

"∴\frac{x}{2}=\frac{0.6}{100}⇒∴x=0.0032"

"K_C=K_P=\frac{x^2}{(1−x)^2}"

"Kp=\frac{(0.0032)^2}{(1−0.0032)^2}=1.03×10^{−5}"